Year           Cash Flow

0 -114,000

1                    48,000

2                   71,000

3                  104,000

Rate = 12%

Calculate for

1.)        Payback Period

2.)        The Internal Rate of Return

3.)        The NPV at capital cost

4.)        The Modified Internal Rate of Return
Answer #1

{Part A:

Payback period is the period in which initial investment is recoevred.

Year Opening Bal CF Closing Bal
1 $ 1,14,000.00 $    48,000.00 $      66,000.00
2 $    66,000.00 $    71,000.00 $      -5,000.00
3 $     -5,000.00 $ 1,04,000.00 $ -1,09,000.00

PBP = Year in which least +ve CB + [ CB in that Year / CF in Next yearc ]

= 1 + [ 66000 / 71000 ]

= 1 + 0.93

= 1.93 Years

Part B:

IRR is the Rate at which PV of Cash Inflows are equal to PV of Cash outflows

Year CF PVF @36% Disc CF PVF @37% Disc CF
0 $ -1,14,000.00     1.0000 $ -1,14,000.00     1.0000 $ -1,14,000.00
1 $      48,000.00     0.7353 $      35,294.12     0.7299 $      35,036.50
2 $      71,000.00     0.5407 $      38,386.68     0.5328 $      37,828.33
3 $ 1,04,000.00     0.3975 $      41,344.39     0.3889 $      40,445.63
NPV $        1,025.19 $          -689.54

IRR = Rate at which least +ve NPV + [ NPV at that Rate / Change in NPV due to 1% inc in DIsc Rate ] * 1%

= 36% + [ 1025.19 / 1714.72 ] * 1%

= 36% + 0.60%

= 36.60%

Part C:

NPV = PV of cash Inflows - PV of cash Outflows

Year CF PVF @12% Disc CF
0 $ -1,14,000.00     1.0000 $ -1,14,000.00
1 $      48,000.00     0.8929 $      42,857.14
2 $      71,000.00     0.7972 $      56,600.77
3 $ 1,04,000.00     0.7118 $      74,025.15
NPV $      59,483.05

Part D:

Modified IRR is similar to IRR. here intermediary Cfs are assumed as reinvested at Cost of Capital.

Year CF FVF @12% FV of CF
1 $    48,000.00       1.2544 $    60,211.20
2 $    71,000.00       1.1200 $    79,520.00
3 $ 1,04,000.00       1.0000 $ 1,04,000.00
FV of CFs $2,43,731.20

Thus 114000 has become 243731.2 over 3 years.

FV = PV(!+r)^n

243731.2 = 114000(1+r)^3

(1+r)^3 = 243731.2 / 114000

= 2.14

1+r = 2.14^ ( 1/3)

= 1.2883

r = 1.2883 - 1

= 0.2883 i.e 28.83%